# secReferences in the Online Appendix

### v3

@@ 1 Introduction
Over recent decades, non-governmental organizations (NGOs) have emerged as influential social actors, shaping regulations and corporate strategies outside traditional political channels (Baron, 2001; Egorov and Harstad, 2017). However, their actions are often difficult to observe, and evidence on the incentives they face and how these shape their strategic decisions remains limited (Yaziji and Doh, 2009). Thus, NGOs are frequently perceived as passive actors, either selecting the most favorable action from a common good perspective (Daubanes and Rochet, 2019) or being dysfunctional and susceptible to corporate capture (Bertrand et al., 2020). This paper introduces a novel approach to analyzing how NGOs make decisions with implications for corporations and their stakeholders.
Consider an NGO with information about corporate wrongdoing that faces a strategic choice: launch an immediate campaign to foster dialogue or wait for the optimal moment when the firm is most sensitive to external pressure. While early action encourages engagement and potential change, strategic timing can amplify impact by signaling strong opposition among stakeholders.
A time when firms are particularly sensitive to external pressure is during their Annual General Meetings (AGMs), recurring shareholder meetings where strategic decisions are revealed and executives are held accountable. We show that NGOs are four to six times more likely to campaign on a target’s AGM date than at other times, a pattern driven by the heightened media coverage AGMs attract. This visibility helps NGOs reach stakeholders and the broader public while imposing reputational costs on firms. However, while AGMs provide a platform for NGOs to maximize exposure, their direct influence is limited—most shareholders vote before the meeting via proxy, reducing the immediate impact on voting outcomes (Bebchuk et al., 2020). Empirically, we leverage the exogenous timing of AGMs, which remains stable within firms due to SEC regulations, allowing us to focus on NGOs’ actions while holding firms’ strategic choices constant.
Our findings reveal the “life cycle” of NGO strategies (Wernerfelt, 1988; Cabral, 2000). Early-stage NGOs prioritize visibility by campaigning on AGM dates, despite their limited direct influence. We show that AGM-timed campaigns prompt consumer backlash and increase the likelihood of related shareholder proposals the following year. However, while these campaigns generate exposure, they do not directly drive change, as, for instance, the target’s environmental and social scores remain unchanged. Over time, NGOs targeting AGMs start campaigning before AGMs to maximize direct influence on corporate decisions. Early campaigns increase the likelihood of related shareholder proposals succeeding, ultimately driving measurable improvements in firms’ environmental and social performance.
Our interpretation is that AGM exposure can enhance NGO reputation. Supporting this view, we find that the success of a shareholder proposal on a given cause increases when an NGO campaigns on the same issue and year, with past AGM-date campaigns by the NGO amplifying this effect. However, the returns to AGM-date campaigns diminish as their frequency rises. In contrast, past NGO campaigns on non-AGM dates have no impact.
We view AGM-date campaigns as a way for NGOs to signal their type to stakeholders and build their “reputation,” as evidenced by the diminishing returns of AGM-targeting over time, which is consistent with reputation models (Bar-Isaac and Tadelis, 2008). [^1] This practice also generates additional negative news to a target firm only for an NGO’s first few campaigns. Later campaigns inflict no additional damage. Thus, the NGO may be willing to incur a cost to build its “reputation-for-toughness,” which it can then leverage in future campaigns (Kreps and Wilson, 1982; Milgrom and Roberts, 1982). Hence, more mature NGOs shift to campaign before AGMs. Furthermore, NGO reputation is likely multidimensional, shaped both by a sustained track record of activism and by a demonstrated commitment to a cause. As such, NGOs may transition to early campaigning not only due to diminishing returns on visibility but also to establish credibility through constructive engagement with key stakeholders (Baur and Palazzo, 2011; Mitchell and Stroup, 2017).
@@ 4.1.1 Shareholders
Figure B7 shows that only campaigns timed twelve months prior to the next AGM increase the likelihood of a shareholder proposal on the same cause. This substantial increase is as large as 12 percentage points, or 44%, above the average probability of observing a proposal. Thus, NGO campaigns from the previous AGMs likely lead to a shareholder proposal at the current AGM, potentially due to heightened media attention.
@@ 4.1.2 Consumers
$\frac{\text{Sales}_{n,q}}{\text{Assets}_{n,q}}$ $=\sum^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}\cdot\text{NGO Campaign}_{n}\cdot\mathbf{1}{\{\text{distance}=\tau\}}\right)+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$ (3)
where $q$ indicates a quarter and $\alpha_{n}$ and $\alpha_{s(n),q}$ are fixed effects at either the firm level or the industry-by-quarter-by-year level, respectively. The dependent value normalizes $n$’s sales by its asset to account for variation in relative firm sizes, which is large in our sample (10 years, largest 3,000 firms). Standard errors are clustered at the firm level.
Appendix Figure B8 reports the estimated $\{\hat{\beta}_{\tau}\}_{\tau}$ coefficients separately for two subsets of the data: for NGO campaigns (i) not planned on an AGM date (blue dots) and (ii) those planned on an AGM date (red triangles). We find that consumers do not respond to NGO campaigns not timed on an AGM day, but they do so to campaigns targeted on the AGM date, implying a loss of sales to total assets of 0.7pp, or about 5%—a magnitude consistent with other papers (Conway and Boxell, 2024; Saint-Jean, 2024).
@@ 4.2 Does AGM-Targeting Create Positive Change?
If public shaming were an effective strategy for influencing a firm’s operations, targeting the AGM month should lead to the greatest improvements in a firm’s E&S score. However, Appendix Figure B9 shows that only campaigns launched 3 to 6 months before an AGM result in a subsequent increase in firms’ E&S scores, with a 2 to 3% rise over the following four quarters. In contrast, campaigns launched at other times, including during the AGM month, have an economically and statistically insignificant effect on firms’ scores.
@@ 6.2 NGO Success and Reputation
Finally, as in signaling games where there are decreasing returns to signal one’s type if the agent has already signaled the type in the past, we estimate decreasing returns from launching campaigns on AGM dates. Column (4) replaces “$\text{Experience}_{g,y}$” for the cumulative count of the campaigns when NGO $g$ targeted an AGM date (“$\text{AGM Experience}_{g,y}$”) and its squared value: the last two rows of Column (4) find that the coefficients of the interactions between those variables and “$\text{Campaign on Same Cause}_{i,n,g,y}$” imply a concave relationship between past AGM-targeting and proposals’ success. Therefore, the support of NGO’s to a shareholder proposal display diminishing returns.
@@ Appendix B Omitted Figures
Figure B8: NGO campaigns and consumer reactions
Note: This figure plots the estimated coefficients of the regression of sales over total assets, on a set of time dummies interacted with a dummy indicating the occurrence of an NGO campaign targeting firm $n$. The regressions also include firm and quarter-by-industry fixed effects. The number of observations is 351,351 and 10,188, respectively. Standard errors are clustered at the firm level.
@@
$\Delta_{4}\text{E\&S Score}_{n,q}=\sum_{l=-6}^{5}\text{Campaign}_{n,q,l}+\textbf{X}_{n,q}\gamma^{\prime}+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$

### v4

@@ 1 Introduction
Over recent decades, non-governmental organizations (NGOs) have emerged as influential actors shaping regulations and corporate strategies outside traditional political channels (Baron, 2001; Egorov and Harstad, 2017). However, their actions are difficult to observe, and evidence on their strategic incentives remains limited (Yaziji and Doh, 2009). As a result, NGOs are often viewed either as principled actors promoting social good (Daubanes and Rochet, 2019) or as dysfunctional entities vulnerable to corporate capture (Bertrand et al., 2020).
This paper provides new insights into how NGO campaigns coordinate stakeholder actions and ultimately shape the creation of corporate social value. We focus on campaign timing—one of the few observable aspects of NGO activity, which remains far more opaque than that of for-profit firms. [^1] A parallel literature uses administrative data to study charitable donations to NGOs (e.g., Cagé et al., 2023; Cagé and Guillot, 2025), offering valuable insights into their funding base. Our focus is instead on NGOs’ strategic choices—an important aspect of their behavior that remains relatively underexplored. While timing may seem like a minor feature, our approach reveals a key trade-off NGOs face: gaining media exposure, which can be leveraged in future campaigns, versus exerting direct influence on firm behavior. We identify the main determinants of how NGOs resolve this trade-off and show that their choices shape which stakeholders are mobilized and how firms ultimately respond.
Our findings reveal a “life cycle” in NGO strategies (Wernerfelt, 1988; Cabral, 2000). Early-stage NGOs prioritize visibility by campaigning on a target firm’s Annual General Meeting (AGM) date, where executives face shareholder scrutiny and heightened media attention, triggering consumer backlash and prompting future shareholder proposals. However, these high-visibility campaigns rarely improve firms’ environmental and social outcomes directly. Over time, NGOs shift to campaigning before a target’s AGM date to exert direct influence, increasing the success rate of current shareholder proposals and leading to measurable improvements in corporate behavior.
Consistent with reputation-building models, shareholder proposals are more likely to succeed when accompanied by NGO campaigns on the same issue—especially when the NGO has a history of targeting AGM dates (Bar-Isaac and Tadelis, 2008). However, this effect diminishes with the number of past AGM-date campaigns, suggesting decreasing returns. In contrast, past non-AGM campaigns have no influence on proposal outcomes. NGOs thus strategically use AGM-date campaigns to signal toughness and build their reputation, initially benefiting from substantial media attention but paying the cost of limited direct influence on AGM voting (Kreps and Wilson, 1982; Milgrom and Roberts, 1982). As a result, more mature NGOs shift toward earlier, pre-AGM campaigns, aiming to build credibility through constructive stakeholder engagement (Baur and Palazzo, 2011; Mitchell and Stroup, 2017).
Our identification strategy leverages the timing of firms’ AGMs, which is exogenously fixed by SEC regulations and stable within firms over time. This enables us to analyze how NGOs respond to predictable opportunities for visibility, holding firm strategies constant. Our main endogenous variable, therefore, is NGO campaign timing, which we explain through NGO experience: more experienced NGOs campaign earlier, trading visibility for impact. While unobservable factors might also influence timing decisions, we verify that key characteristics of the target firms—such as size and ESG scores—do not systematically correlate with NGO experience. This reduces concerns about selection bias, for instance, ruling out that less experienced NGOs strategically target smaller or less prosocial firms for easier gains.
@@ 4.1.1 Shareholders
Figure B7 shows that only campaigns launched twelve months before the next AGM increase the likelihood of a shareholder proposal on the same issue. The effect is substantial—an increase of 12 percentage points, or 44% relative to the baseline. Since SEC 14a8(e)(2) rule prevents shareholders from submitting proposals within 120 days of the proxy statement’s release, the x-axis is capped at four months before the previous year’s AGM. This timing suggests that campaigns shortly after an AGM may spur proposals at the next one, possibly due to heightened media attention.
@@ 4.1.2 Consumers
$\frac{\text{Sales}_{n,q}}{\text{Assets}_{n,q-1}}$ $=\sum^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}\cdot\text{NGO Campaign}_{n}\cdot\mathbf{1}{\{\text{distance}=\tau\}}\right)+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$ (3)
where $q$ indicates a quarter and $\alpha_{n}$ and $\alpha_{s(n),q}$ are fixed effects at either the firm level or the industry-by-quarter-by-year level, respectively. The dependent value normalizes $n$’s sales by its asset to account for time variation in relative firm sizes, which is large in our ten-year sample. Standard errors are clustered at the firm level.
Appendix Figure B8 shows the estimated ${\hat{\beta}\tau}\tau$ coefficients separately for NGO campaigns (i) planned on an AGM date (Panel a) and (ii) not planned on an AGM date (Panel b). We find no consumer response to the latter, but sales drop by 0.7pp (about 5%) for the former—a magnitude consistent with Conway and Boxell (2024); Saint-Jean (2024). In Panel (c), the sales gap between AGM and non-AGM campaigns emerges one quarter before the event and continues to widen, possibly due to the greater media coverage of AGM dates.
@@ 4.2 Does AGM-Targeting Create Positive Change?
If public shaming were an effective strategy to influence firm behavior, targeting the AGM month should yield the largest improvements in E&S scores. However, Appendix Figure B9 shows that only campaigns launched 3 to 6 months before the AGM lead to subsequent increases in firms’ E&S scores—rising by 2–3% over the following year. By contrast, campaigns launched at other times, including during the AGM month, have no economically or statistically significant effect.
@@ 6.2 NGO Success and Reputation
Finally, as in signaling games where there are decreasing returns to signal one’s type if the agent has already signaled the type in the past, we estimate decreasing returns from launching campaigns on AGM dates. Column (4) replaces “$\text{Experience}_{g,y}$” for the cumulative count of the campaigns when NGO $g$ targeted an AGM date (“$\text{AGM Experience}_{g,y}$”) and its squared value: the last two rows of Column (4) find that the coefficients of the interactions between those variables and “$\text{Campaign on Same Cause}_{i,n,g,y}$” imply a concave relationship between past AGM-targeting and proposals’ success. Therefore, the support of NGO’s to a shareholder proposal display diminishing returns. [^13] Note that the combined effect of AGM Experience and AGM Experience Squared never yields a negative probability of proposal success—even at the highest observed levels of AGM Experience—suggesting our reputation proxy does not exhibit negative returns.
@@ Appendix B Omitted Figures
Figure B8: Examples of positive and negative E&S news
(a) NGO Campaigns on AGM
(b) NGO Campaigns on Other Days
(c) Estimated Effect of Campaigning on AGM
Notes: Panels (a) and (b) plot the estimated coefficients of the regression of sales over (lagged) total assets, on a set of time dummies interacted with a dummy indicating the occurrence of an NGO campaign targeting firm $n$ from (3). Panel(c) plots the estimated $\hat{\beta}$ from
$\frac{\text{Sales}_{n,q}}{\text{Assets}_{n,q-1}}$ $=\sum^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}\cdot\text{NGO Campaign on AGM}_{n}\cdot\mathbf{1}{\{\text{distance}=\tau\}}\right)$
$+\sum^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\tilde{\beta}_{\tau}\cdot\mathbf{1}{\{\text{distance}=\tau\}}\right)+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$
which interacts the time dummy, the occurrence of a campaign, and a dummy indicating that the campaign happened on AGM day. The regressions also include firm and quarter-by-industry fixed effects. The reference period is -1 quarter, consistent with the other analyses. The number of observations is 10,188, 351,351, and 361,539, respectively. Standard errors are clustered at the firm level.
@@
$\Delta_{4}\text{E\&S Score}_{n,q}=\sum_{l=-6}^{5}\text{Campaign}_{n,q,l}+\textbf{X}_{n,q}\gamma^{\prime}+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$ (B1)
@@
Unlike an event study, equation (B1) is a distributed leads-and-lags regression. We omit a reference period, so the coefficients are interpreted relative to firms that did not experience a campaign. This choice allows us to directly compare the effects of early campaigns with those launched on the AGM date relative to firms without campaigns. Results remain similar if we instead set month –4 as the reference period, as in the other analyses.